# Discovery of Seven Cold and Distant Brown Dwarfs with JWST RUBIES

### v1

@@ Abstract
We report near-infrared spectral model fits to seven distant L- and T-type dwarfs observed with the JWST Near Infrared Spectrograph (NIRSpec) as part of the Red Unknowns: Bright Infrared Extragalactic Survey (RUBIES). Comparison of 0.9–2.5 m near-infrared spectra of these sources to spectral standards indicates spectral types spanning L1 to T8, and spectrophotometric distances spanning 800–3,000 pc. Fits to three grids of spectral models yield atmosphere parameters and spectrophotometric distances largely consistent with our classifications, although fits to L dwarf spectra indicate missing components to the models. Three of our sources have vertical displacements from the Galactic plane exceeding 1 kpc, and have high probabilities of membership in the Galactic thick disk population. Of these, the L dwarf RUBIES-BD-3 (RUBIES-EGS-3081) is well-matched to subdwarf standards, while the early T dwarf RUBIES-BD-5 (RUBIES-UDS-170428) is best fit by metal-poor atmosphere models; both many be a thick disk or halo brown dwarfs. We critically examine the 1–5 m spectra of the current sample of 1–2 kpc mid- and late-T dwarfs, finding that temperature, surface gravity, metallicity, and vertical mixing efficiency can all contribute to observed variations in near-infrared spectral structure and the strength of the 4.2 m CO2 band. This work aims to guide ongoing JWST, Euclid, and other space-based spectral surveys that are expected to uncover thousands of low-temperature stars and brown dwarfs throughout the Milky Way.
@@ 1 Introduction
Brown dwarfs are the lowest-temperature and lowest-mass “stars” in the Milky Way, with insufficient mass (0.075 M⊙) to sustain core hydrogen fusion (Kumar, 1962, 1963; Hayashi & Nakano, 1963). While comprising a significant fraction of star-like objects in the Galactic disk (20%; Reylé 2018; Kirkpatrick et al. 2021), these objects cool and dim over time, becoming exceedingly faint ( 18, 10; Dahn et al. 2017). As such, the majority of low-temperature stars and brown dwarfs—classified as late-M, L, T and Y dwarfs—are detected in the immediate solar neighborhood (100 pc), and have predominantly solar elemental compositions. Only a small fraction of local dwarfs are metal-poor members of the Galactic thick disk (12%) and halo populations (0.5%; Jurić et al. 2008; Schneider et al. 2020). While rare, these long-lived sources are essential for understanding how metallicity influences the formation, evolution, and hydrogen-burning mass threshold for brown dwarfs, and can potentially trace dynamical evolution and chemical enrichment throughout the history of the Milky Way (Burgasser, 2004a; Honaker & Gizis, 2025).
@@
The high infrared sensitivity of the current generation of space-based telescopes, including JWST (Rigby et al., 2023) and Euclid (Laureijs et al., 2011), has made it possible to identify and study brown dwarfs at considerably larger distances. Building on tools developed for deep imaging and spectroscopic surveys conducted with the Hubble Space Telescope (HST; Pirzkal et al. 2009; Ryan et al. 2011; van Vledder et al. 2016; Aganze et al. 2022a), deep JWST surveys are now uncovering individual low-temperature stars and brown dwarfs beyond kiloparsec distances (Nonino et al., 2023; Langeroodi & Hjorth, 2023; Wang et al., 2023; Burgasser et al., 2024; Holwerda et al., 2024; Hainline et al., 2024a, b; Tu et al., 2025a, b; Hainline et al., 2025). Early Euclid studies are uncovering hundreds of cool dwarfs to similar scales (Žerjal et al., 2025; Mohandasan et al., 2025; Dominguez-Tagle et al., 2025). Those surveys conducted at high galactic latitudes contain considerably higher fractions of thick disk and halo sources than in the local population, enabling more efficient and comprehensive studies of metal-poor low-mass stars and brown dwarfs.
@@
In this article, we present the identification and spectral analysis of seven L and T dwarfs identified in deep spectroscopic data collected as part of the Red Unknowns: Bright Infrared Extragalactic Survey (RUBIES; de Graaff et al. 2025). In Section 2, we describe the design and data associated with the RUBIES survey, and our method for identifying brown dwarfs from the spectral data. In Section 3, we present an empirical analysis of the spectra, including template classification and spectrophotometric distance estimates. In Section 4, we fit the spectra to three sets of model grids to obtain robust estimates of atmospheric parameters and their uncertainties, and evaluate the accuracy of the different grids over the broad range of specral types in our sample. In Section 5, we quantify the probabilities of membership in the Galactic thin disk, thick disk, and halo populations, and critically assess the source of spectral variations among mid- and late-type T dwarfs in our sample. We summarize our results in Section 6.
@@ 2.1 RUBIES Survey Data
RUBIES is a spectroscopic survey conducted with the JWST Near Infrared Spectrograph (NIRSpec; Birkmann et al. 2022). The survey is focused on detecting and characterizing massive galaxies at high redshifts, drawing on sources selected from JWST Near Infrared Camera (NIRCam) imaging data obtained in the Extended Groth Strip (EGS) by the Cosmic Evolution Early Release Science Survey (CEERS, JWST-ERS-1345, PI S. Finkelstein), and in the Ultra Deep Survey (UDS) by the Public Release IMaging for Extragalactic Research Survey (PRIMER, JWST-GO-1837, PI J. Dunlap). RUBIES prioritizes very red sources, and thus has a higher chance of observing brown dwarfs than surveys with differing or without color selection. All RUBIES sources were acquired using the NIRSpec micro-shutter array (MSA; Ferruit et al. 2022) with a three-shutter slitlet, with total integrations of 48 minutes acquired in each of the low-resolution Prism/Clear mode (0.6–5.3 m, = 30–300) and the medium-resolution G395M/F290LP mode (2.9–5.3 m, = 1000), in both cases using a three-point nodding pattern. Spectra were reduced using the JWST calibration pipeline version 1.12.5 (Bushouse et al., 2023) and msaexp (Brammer, 2023) as described in detail in Heintz et al. (2025) and de Graaff et al. (2025). Here, we focus on the NIRSpec/Prism data which provide the highest signal-to-noise (S/N) for our faint targets.
@@ 2.2 Selection of Brown Dwarfs
| Other ID | o006_s00089aaAs reported in Tu et al. (2025a). | o006_s35616aaAs reported in Tu et al. (2025a). | o005_s41280aaAs reported in Tu et al. (2025a). | ||||
| CEERS-EGS-BD-4bbAs reported in Hainline et al. (2024a). | |||||||
@@
Our seven sources were selected among point sources identified in the CEERS/EGS and PRIMER/UDS NIRCam imaging, with spectral data grades 2.5 (a quality assessment metric indicating the reliability of spectral features) and automated redshift estimates of . These criteria initially narrowed our selection to six sources in the EGS field and three sources in the UDS field. We visually inspected NIRCam images to ensure the sources were consistent with point sources, and visually confirmed that the NIRSpec/Prism spectra were consistent with low-temperature dwarf morphologies as opposed to active galactic nuclei or other compact extragalactic sources. This evaluation allowed us to reject two duplicate EGS sources, resulting in a final sample of seven low-temperature stars and brown dwarfs, summarized in Table 1. We list total flux photometry in the NIRCam F115W, F150W, F200W, F277W, F356W, and F444W filters from the CEERS/EGS and PRIMER/UDS surveys as compiled in the ASTRODEEP-JWST catalog (Merlin et al., 2024). For the corresponding magnitudes, we conservatively include a systematic uncertainty of 0.05 mag to account for zeropoint errors. Three of the sources—RUBIES-BD-1, -2, and -3—were previously reported by Hainline et al. (2024a) and Tu et al. (2025a), whose results we discuss below in comparison with our findings.
@@
Figure 1 displays the NIRSpec/Prism spectra for these sources, flux-calibrated to their apparent F444W magnitudes and presented in flux density units. RUBIES-BD-1, -6, and -7 exhibit sharp flux peaks at 1.0, 1.25, 1.6, and 2.1 m, and a broad peak at 4 m, shaped by strong H2O and CH4 bands that are typical of T dwarfs (Burgasser et al., 2006; Sorahana & Yamamura, 2012; Beiler et al., 2024a). RUBIES-BD-2 displays broader flux peaks at 1.2, 1.65, and 2.15 m between strong H2O absorption bands; and a blue, relatively featureless spectral slope toward longer wavelengths, typical of L dwarfs (Cushing et al., 2005; Sorahana & Yamamura, 2012). RUBIES-BD-3, -4, and -5 have features intermediate between these morphological types (only marginally discernible in the spectrum of our faintest target RUBIES-BD-4), suggesting that they are late-L or early-T dwarfs. RUBIES-BD-3 also shows a peculiar split in the 1.2 m band and a bluer overall spectral morphology which may be associated with subsolar metallicity, as discussed below.
@@ 2.3 Classification and Distance Estimation
Spectral classifications for these sources were determined by comparing their 0.9–2.4 m spectra to L and T dwarf spectral standards defined in Kirkpatrick et al. (2010) and Burgasser et al. (2006), respectively. Near-infrared spectral data for these standards were drawn from the SpeX Prism Library Analysis Toolkit (Burgasser & Splat Development Team, 2017); see associated references in Table 7 of the Appendix. We also selected low-metallicity L and T dwarf spectral standards encompassing mild subdwarf (d/sd), subdwarf (sd), and extreme subdwarf (esd) classes as defined in Zhang et al. (2017b, 2018a); Greco et al. (2019) and Burgasser et al. (2025b), drawing on published near-infrared spectral data acquired with various instruments for comparison. All spectra were smoothed and interpolated onto the same wavelength grid corresponding to the NIRSpec/prism data. For our source spectra, we also applied small fractional pixel shifts of up to 1.5 pixels before interpolation, as determined by visual comparison to the best-fit standards, to account for variations in the projection of spectral data on the NIRSpec detector from the MSA slits.
@@
(cf. Cushing et al. 2005). For RUBIES-BD-4 and -5, the spectral S/N is sufficiently low that the statistic cannot differentiate between adjacent standards, so we conducted a “by-eye” comparison to the spectral standards to identify an approximate best fit while allowing for generous uncertainties.
@@
Sources and their associated best-fit standards are shown in Figure 2 and summarized in Table 2. RUBIES-BD-1, -2, -6, and -7 are well-matched to T7, L1, T8, and T5 dwarf standards, respectively, and we estimate uncertainties of 0.5 subtypes for these sources. For RUBIES-BD-4 and -5, our best “by-eye” fits are to the T2 and T1 standards, respectively, with deviations generally consistent with the lower S/N of the spectra. For these sources, we estimate larger classification uncertainties of 3 and 2 subtypes, respectively. For RUBIES-BD-3, the minimum among the dwarf standards is achieved for the L3 dwarf, but there are significant deviations in the 1.0–1.4 m -band region. We therefore compared this spectrum to the broader subdwarf grid (Figure 3), finding the d/sdL7 mild subdwarf and sdL5 subdwarf standards to be the best fits (minimum ). We adopt a classification of sd:L61 for this source.
With these classifications, we estimated distances to each source based on the NIRCam photometry, using the spectral type/absolute magnitude conversions from Burgasser et al. (2024) based on the Sonora Bobcat models (Marley et al., 2021). For RUBIES-BD-1, -2, -4, -5, -6, and -7 we assumed a metallicity [M/H] = 0.0; for RUBIES-BD-3 we assumed [M/H] = 0.5. Uncertainties include contributions from the measured photometry and spectral classifications, and were combined using Monte Carlo techniques. Distance estimates are listed in Table 2, and are in agreement across all six filters for all of the sources with the exception of RUBIES-BD-3, which varies from pc for F115W photometry to pc for F356W photometry. This 3.6 discrepancy is likely associated with the metal-poor nature of the source. We computed uncertainty-weighted average distance estimates by using the higher of the lower and upper error bounds as the weighting factor (), and computed the combined uncertainty as
where = 3–6 is the number of filters with distance estimates. This expression accounts for both the scatter among distance estimates and individual distance uncertainties. These distances range from 83090 pc for RUBIES-BD-7 to nearly 3 kpc for RUBIES-BD-4, with five of the sources having estimated distances exceeding 1 kpc. At the Galactic latitude of the EGS and UDS fields, these distances correspond to vertical offsets from the disk ( = ) spanning of 720 to 2600 pc, extending into the volume dominated by thick disk objects (Jurić et al., 2008; Burgasser et al., 2024).
@@ 3.1 Model Grids
To characterize the physical properties of these sources, we compared the NIRSpec spectra to three different model sets, allowing us to assess the quality of the fits across different model assumptions. Table 3 summarizes the models and associated parameter ranges explored. The Sonora Elf Owl models (Mukherjee et al., 2024) are cloud-free radiative-convective equilibrium models that include the effects of disequilibrium chemistry due to vertical mixing, and allow for variations in temperature (Teff), surface gravity (), bulk metallicity ([M/H]), the C/O elemental abundance ratio, and the vertical diffusion coefficient (). We use the version of these models that have artificially suppressed PH3 abundances (see discussion in Beiler et al. 2024b). The Sonora Diamondback models (Morley et al., 2024) are based on the same framework but include refractory clouds in both the chemistry and atmospheric opacity using the approach of Ackerman & Marley (2001). These models are computed in chemical equilibrium and include an parameter to account for the efficiency of condensate sedimentation. Here, we adopt a value of = 10 to denote models without clouds. The SAND (Spectral ANalog of Dwarfs) model set (Gerasimov et al., 2024; Alvarado et al., 2024) is based on the PHOENIX code (Hauschildt et al., 1997), and was designed to explore the non-solar abundances of brown dwarfs in globular clusters and the Milky Way’s thick disk and halo. This grid samples variations in alpha element enrichment ([/Fe]), accounts for self-consistent condensation and gravitational settling of clouds following Allard et al. (2012), and assumes chemical equilibrium.
@@
| C/ObbRelative the Sun, for which we adopt a value of C/O = 0.458 (Lodders, 2003). | 0.5 to 1.5 | 0.05 | |||
| ccA value of = 10 is adopted for Diamondback models without clouds; see Morley et al. (2024). | 1 to 10 | 0.25 | |||
@@
All data and models were interpolated onto a common wavelength scale that matches the variable resolution and wavelength sampling of the NIRSpec/Prism mode, and were evaluated in flux density units scaled to each source’s apparent F444W magnitudes. Because the models are computed in surface flux densities, the optimal scaling factory between data and models (Eqn. 2) is equal to the ratio , where is the radius of the source and its distance. We assumed a fixed radius of 0.8 RJup, expected for an evolved, relatively massive brown dwarf (Marley et al., 2021), so that the scale factor provides a second estimate of the source distance of = 1.8510 pc, which can be compared to the distances inferred from spectrophotometry.

### v2

@@ Abstract
We report near-infrared spectral model fits to seven distant L- and T-type dwarfs observed with the JWST Near Infrared Spectrograph (NIRSpec) as part of the Red Unknowns: Bright Infrared Extragalactic Survey (RUBIES). Comparison of 0.9–2.5 m near-infrared spectra of these sources to spectral standards indicates spectral types spanning L1 to T8 and spectrophotometric distances spanning 800–3,000 pc. Fits to three grids of spectral models yield atmosphere parameters and spectrophotometric distances largely consistent with our classifications, although fits to L dwarf spectra indicate missing components to the models. Three of our sources have vertical displacements from the Galactic plane exceeding 1 kpc, and have high probabilities of membership in the Galactic thick disk population. Of these, the L dwarf RUBIES-BD-3 (RUBIES-EGS-3081) is well-matched to subdwarf standards, while the early T dwarf RUBIES-BD-5 (RUBIES-UDS-170428) is best fit by metal-poor atmosphere models; both be a thick disk or halo brown dwarfs. We critically examine the 1–5 m spectra of the current sample of 1–2 kpc mid- and late-T dwarfs, finding that temperature, surface gravity, metallicity, and vertical mixing efficiency can all contribute to observed variations in near-infrared spectral structure and the strength of the 4.2 m CO2 band. This work aims to guide ongoing JWST, Euclid, and other space-based spectral surveys that are expected to uncover thousands of low-temperature stars and brown dwarfs throughout the Milky Way.
@@ 1 Introduction
Brown dwarfs are the lowest-temperature and lowest-mass “stars” in the Milky Way, with insufficient mass (0.075 M⊙) to sustain core hydrogen fusion (1962AJ.....67S.579K; 1963ApJ...137.1121K; 1963PThPh..30..460H). While comprising a significant fraction of star-like objects in the Galactic disk (20%; Reyl__2018; Kirkpatrick_2021), these objects cool and dim over time, becoming exceedingly faint ( 18, 10; 2017AJ....154..147D). As such, the majority of low-temperature stars and brown dwarfs—classified as late-M, L, T and Y dwarfs—are detected in the immediate solar neighborhood (100 pc), and have predominantly solar elemental compositions. Only a small fraction of local dwarfs are metal-poor members of the Galactic thick disk (12%) and halo populations (0.5%; 2008ApJ...673..864J; 2020ApJ...898...77S). While rare, these long-lived sources are essential for understanding how metallicity influences the formation, evolution, and hydrogen-burning mass threshold brown dwarfs, and can potentially trace dynamical evolution and chemical enrichment throughout the history of the Milky Way (2004ApJS..155..191B; 2025ApJ...985...48H).
@@
The high infrared sensitivity of the current generation of space-based telescopes, including JWST (2023PASP..135d8001R) and Euclid (2011arXiv1110.3193L), has made it possible to identify and study brown dwarfs at considerably larger distances. Building on tools developed for deep imaging and spectroscopic surveys conducted with the Hubble Space Telescope (2009ApJ...695.1591P; 2011ApJ...739...83R; 2016MNRAS.458..425V; 2022ApJ...924..114A), deep JWST surveys are now uncovering individual low-temperature stars and brown dwarfs beyond kiloparsec distances (2023ApJ...942L..29N; 2023ApJ...957L..27L; 2023MNRAS.523.4534W; 2024ApJ...962..177B; 2024MNRAS.529.1067H; 2024ApJ...964...66H; 2024ApJ...975...31H; 2025ApJ...980..230T; 2025arXiv251002026T; 2025arXiv251000111H). Early Euclid studies are uncovering hundreds of cool dwarfs to similar scales (2025arXiv250322497V; 2025arXiv250322559M; 2025ApJ...991...84D). Those surveys conducted at high galactic latitudes contain considerably higher fractions of thick disk and halo sources than in the local population, enabling more efficient and comprehensive studies of metal-poor low-mass stars and brown dwarfs.
@@
In this article, we present the identification and spectral analysis of seven L and T dwarfs identified in deep spectroscopic data collected as part of the Red Unknowns: Bright Infrared Extragalactic Survey (RUBIES; 2025A&A...697A.189D). In Section 2, we describe the design and data associated with the RUBIES survey, and our method for identifying brown dwarfs from the spectral data. In Section 3, we present an empirical analysis of the spectra, including template classification and spectrophotometric distance estimates. In Section 4, we fit the spectra to three sets of model grids to obtain robust estimates of atmospheric parameters and their uncertainties, and evaluate the accuracy of the different grids over the broad range of specral types in our sample. In Section 5, we quantify the probabilities of membership in the Galactic thin disk, thick disk, and halo populations, and critically assess the source of spectral variations among mid- and late-type T dwarfs in our sample. We summarize our results in Section 6.
@@ 2.1 RUBIES Survey Data
RUBIES is a spectroscopic survey conducted with the JWST Near Infrared Spectrograph (NIRSpec; Birkmann_2022). The survey is focused on detecting and characterizing massive galaxies at high redshifts, drawing on sources selected from JWST Near Infrared Camera (NIRCam) imaging data obtained in the Extended Groth Strip (EGS) by the Cosmic Evolution Early Release Science Survey (CEERS, JWST-ERS-1345, PI S. Finkelstein), and in the Ultra Deep Survey (UDS) by the Public Release IMaging for Extragalactic Research Survey (PRIMER, JWST-GO-1837, PI J. Dunlap). RUBIES prioritizes very red sources, and thus has a higher chance of observing brown dwarfs than surveys with differing color selection. All RUBIES sources were acquired using the NIRSpec micro-shutter array (MSA; 2022A&A...661A..81F) with a three-shutter slitlet, with total integrations of 48 minutes acquired in each of the low-resolution Prism/Clear mode (0.6–5.3 m, = 30–300) and the medium-resolution G395M/F290LP mode (2.9–5.3 m, = 1000), using a three-point nodding pattern. Spectra were reduced using the JWST calibration pipeline version 1.12.5 (bushouse_2023_10022973) and msaexp (brammer_2023_8319596) as described in detail in 2025A&A...693A..60H and 2025A&A...697A.189D. Here, we focus on the NIRSpec/Prism data which provide the highest signal-to-noise (S/N) for our faint targets.
@@ 2.2 Selection of Brown Dwarfs
| Other ID | o006_s00089aaAs reported in 2025ApJ...980..230T. | o006_s35616aaAs reported in 2025ApJ...980..230T. | o005_s41280aaAs reported in 2025ApJ...980..230T. | ||||
| CEERS-EGS-BD-4bbAs reported in 2024ApJ...964...66H. | |||||||
@@
Our seven sources were selected among point sources identified in the CEERS/EGS and PRIMER/UDS NIRCam imaging , with spectral data grades 2.5 (a quality assessment metric indicating the reliability of spectral features) and automated redshift estimates of . These criteria initially narrowed our selection to six sources in the EGS field and three sources in the UDS field. We visually inspected NIRCam images to ensure the sources were consistent with point sources, and visually confirmed that the NIRSpec/Prism spectra were consistent with low-temperature dwarf morphologies as opposed to active galactic nuclei or other compact extragalactic sources. This evaluation allowed us to reject two duplicate EGS sources111, resulting in a final sample of seven low-temperature stars and brown dwarfs, summarized in Table 1. We list in the NIRCam F115W, F150W, F200W, F277W, F356W, and F444W filters from the CEERS/EGS and PRIMER/UDS surveys as compiled in the ASTRODEEP-JWST catalog (2024A&A...691A.240M). For the corresponding magnitudes, we conservatively include a systematic uncertainty of 0.05 mag to account for zeropoint errors. Three of the sources—RUBIES-BD-1, -2, and -3—were by 2024ApJ...964...66H and 2025ApJ...980..230T, whose results we discuss below in comparison with our findings.
@@
Figure 1 displays the NIRSpec/Prism spectra for these sources, flux-calibrated to their apparent F444W magnitudes and presented in flux density units.222 RUBIES-BD-1, -6, and -7 exhibit sharp flux peaks at 1.0, 1.25, 1.6, and 2.1 m, and a broad peak at 4 m, shaped by strong H2O and CH4 bands that are typical of T dwarfs (Burgasser_2006; 2012ApJ...760..151S; 2024ApJ...973..107B). RUBIES-BD-2 displays broader flux peaks at 1.2, 1.65, and 2.15 m between strong H2O absorption bands; and a blue, relatively featureless spectral slope toward longer wavelengths, typical of L dwarfs (2005ApJ...623.1115C; 2012ApJ...760..151S). RUBIES-BD-3, -4, and -5 have features intermediate between these morphological types (only marginally discernible in the spectrum of our faintest target RUBIES-BD-4), suggesting that they are late-L or early-T dwarfs. RUBIES-BD-3 also shows a peculiar split in the 1.2 m band and a bluer overall spectral morphology which may be associated with subsolar metallicity, as discussed below.
@@ 2.3 Classification and Distance Estimation
Spectral classifications for these sources were determined by comparing their 0.9–2.4 m spectra to L and T dwarf spectral standards defined in 2010ApJS..190..100K and Burgasser_2006, respectively. Near-infrared spectral data for these standards were drawn from the SpeX Prism Library Analysis Toolkit (2017ASInC..14....7B); see associated references in Table 7 of the Appendix. We also selected low-metallicity L and T dwarf spectral standards encompassing mild subdwarf (d/sd), subdwarf (sd), and extreme subdwarf (esd) classes as defined in 2017MNRAS.464.3040Z; 2018MNRAS.479.1383Z; 2019AJ....158..182G; and 2025ApJ...982...79B, drawing on published near-infrared spectral data acquired with various instruments for comparison . All spectra were smoothed and interpolated onto the same wavelength grid corresponding to the NIRSpec/Prism data. For our source spectra, we also applied small fractional pixel shifts of up to 1.5 pixels before interpolation, as determined by visual comparison to the best-fit standards, to account for variations in the projection of spectral data on the NIRSpec detector from the MSA slits.
@@
(cf. 2005ApJ...623.1115C). For RUBIES-BD-4 and -5, the spectral S/N is sufficiently low that the statistic cannot differentiate between adjacent standards, so we conducted a “by-eye” comparison to the spectral standards to identify an approximate best fit while allowing for generous uncertainties.
@@
Sources and their associated best-fit standards are shown in Figure 2 and summarized in Table 2. RUBIES-BD-1, -2, -6, and -7 are well-matched to T7, L1, T8, and T5 dwarf standards, respectively, and we estimate uncertainties of 0.5 subtypes for these sources. For RUBIES-BD-4 and -5, our best “by-eye” fits are to the T2 and T1 standards, respectively, with deviations generally consistent with the lower S/N of the spectra. For these sources, we estimate larger classification uncertainties of 3 and 2 subtypes, respectively. For RUBIES-BD-3, the minimum among the dwarf standards is achieved for the L3 dwarf, but there are significant in the 1.0–1.4 m -band region. We therefore compared this spectrum to the broader subdwarf grid (Figure 3), finding the d/sdL7 mild subdwarf and sdL5 subdwarf standards to be the best fits (minimum ). We adopt a classification of sd:L61 for this source.
With these classifications, we estimated distances to each source NIRCam photometry, using the spectral type/absolute magnitude conversions from 2024ApJ...962..177B based on the Sonora Bobcat models (2021ApJ...920...85M). For RUBIES-BD-1, -2, -4, -5, -6, and -7 we assumed a metallicity [M/H] = 0.0; for RUBIES-BD-3 we assumed [M/H] = 0.5. Uncertainties include contributions from the measured photometry and spectral classifications, and were combined using Monte Carlo techniques. Distance estimates are listed in Table 2, and are in agreement across all six filters for all of the sources with the exception of RUBIES-BD-3, which varies from pc for F115W photometry to pc for F356W photometry. This 3.6 discrepancy is likely associated with the metal-poor nature of the source. We computed uncertainty-weighted average distance estimates by using the higher of the lower and upper error bounds as the weighting factor (), and computed the combined uncertainty as
where = 3–6 is the number of filters with distance estimates. This expression accounts for both the scatter among distance estimates and individual distance uncertainties. Distances range from 83090 pc for RUBIES-BD-7 to nearly 3 kpc for RUBIES-BD-4, with five of the sources having estimated distances exceeding 1 kpc. At the Galactic latitude of the EGS and UDS fields, these distances correspond to vertical offsets from the disk , extending into the volume dominated by thick disk objects (2008ApJ...673..864J; 2024ApJ...962..177B).
@@ 3.1 Model Grids
To characterize the physical properties of these sources, we compared the NIRSpec spectra to three different model sets, allowing us to assess the quality of the fits across different model assumptions. Table 3 summarizes the models and associated parameter ranges explored. The Sonora Elf Owl models (2024ApJ...963...73M) are cloud-free radiative-convective equilibrium models that include the effects of disequilibrium chemistry due to vertical mixing, and allow for variations in temperature (Teff), surface gravity (), bulk metallicity ([M/H]), the C/O elemental abundance ratio, and the vertical diffusion coefficient (). We use the version of these models that have artificially suppressed PH3 abundances (see 2024ApJ...973...60B ). The Sonora Diamondback models (Morley_2024) are based on the same framework but include refractory clouds in both the chemistry and atmospheric opacity using the approach of 2001ApJ...556..872A. These models are computed in chemical equilibrium and include an parameter to account for the efficiency of condensate sedimentation. Here, we adopt a value of = 10 to denote models without clouds. The SAND (Spectral ANalog of Dwarfs) model set (2024ApJ...971...65G; 2024RNAAS...8..134A) is based on the PHOENIX code (1997ApJ...483..390H), and was designed to explore the non-solar abundances of brown dwarfs in globular clusters and the Milky Way’s thick disk and halo. This grid samples variations in alpha element enrichment ([/Fe]), accounts for self-consistent condensation and gravitational settling of clouds following 2012RSPTA.370.2765A, and assumes chemical equilibrium.
@@
| C/ObbRelative the Sun, for which we adopt a value of C/O = 0.458 (2003ApJ...591.1220L). | 0.5 to 1.5 | 0.05 | |||
| ccA value of = 10 is adopted for Diamondback models without clouds; see Morley_2024. | 1 to 10 | 0.25 | |||
@@
References. — [1] 2024ApJ...963...73M; [2] Morley_2024; [3] 2024RNAAS...8..134A; [4] 2024ApJ...971...65G.
All data and models were interpolated onto a common wavelength scale that matches the variable resolution and wavelength sampling of the NIRSpec/Prism mode, and were evaluated in flux density units scaled to each source’s apparent F444W magnitudes. Because the models are computed in surface flux densities, the optimal scaling factory between data and models (Eqn. 2) is equal to the ratio , where is the radius of the source and its distance. We assumed a fixed radius of 0.8 RJup, expected for an evolved, relatively massive brown dwarf.333 (2021ApJ...920...85M) scale factor provides a second estimate of the source distance = 1.8510 pc.